# XEQ Scale for Evaluating XAI Experience Quality Grounded in Psychometric Theory

### v2

@@ 5.1 Study Design and Applications
Figure 1: The static previews of the relatively positive (left 2 columns) and negative (right column) XAI Experiences with the CourseAssist Chatbot designed for the Pilot Study
5.2 Recruitment Details
This study enlisted 203 participants, comprising undergraduate students from the leading research institute and participants recruited from the Prolific.co platform. 33 students from the research institute and 70 Prolific participants were recruited for the CourseAssist application where the inclusion criteria were: Current education level - Undergraduate degree; Degree subjects - Mathematics and statistics, Information and Communication Technologies, Natural sciences; and Year of study - 1st, 2nd, 3rd or 4th. 100 Prolific participants were recruited for the AssistHub application with the following inclusion criteria: Household size is 3 or larger; Property ownership is either social housing or affordable-rented accommodation; Employment status is either part-time, due to start a new job within the next month, unemployed, or not in paid work (e.g. homemaker or retired).
@@ 5.2 Recruitment Details
In the rest of this paper, we will refer to all responses to positive experiences as Group A and all responses to negative experiences as Group B. To represent application-specific groups we will use the application name as a prefix; e.g. CourseAssist-A. Each participant was randomly assigned to one of the sample experiences and after review, we excluded 5, 1 and 1 responses from groups CourseAssist-A, AssistHub-A and AssistHub-B who failed the following attention checks: 1) spend less than half of the allocated time; and/or 2) responded to the questionnaire in a pattern. This resulted in 53, 50, 50, and 50 respondents for CourseAssist-A, CourseAssist-B, AssistHub-A and AssistHub-B groups respectively.
5.3 Metrics
@@ 5.3.3 Construct Validity
Construct validity evaluates the degree to which the scale assesses the characteristic of interest [14]. We perform two forms of Factor Analysis (FA) to uncover underlying factors (i.e. dimensions) and validate them.
Exploratory FA
@@
Confirmatory FA
hypothesise a factor model (e.g. model proposed in the Exploratory FA or a Factor model proposed by XAI experts) and calculate factor loadings where the expected loading for each item is expected to be $\geq 0.5$.
@@ 5.4.1 Internal Consistency
Table 2 column 3 reports the Item-Total Correlation. All items met the baseline criteria of $iT\geq 0.5$ and baseline criteria for Inter-Item correlation. Cronbach’s alpha is $0.9562$ which also indicates strong internal consistency.
@@ 5.4.2 Discriminant Validity
We performed discriminant analysis over 100 trials where at each trial a different train-test split of the responses was used. Each trial used a stratified split, with 70% of the responses for training and 30% for testing. Over the 100 trials, we observed accuracy of $0.63\pm 0.05$ and a macro F1-score of $0.63\pm 0.05$ which is significantly over the baseline accuracy of $0.50$ for a binary classification task. Mixed-model ANOVA test showed a statistically significant difference between groups A and B with a p-value of $121.63e-$ 12 where the mean participant total for groups A and B were $70.96\pm 0.47$ and $57.97\pm 1.84$. Also, it revealed a substantial variability within groups indicated by the group variance of 104.86, which we account for including responses from two application domains. Furthermore, Cohen’s d was 1.7639 which indicates a large effect size confirming a significant difference between groups A and B. A standard t-test also obtained close to zero p-value associated with the t-statistic at $091.13e-$ 09 further verifying statistical difference. Based on this evidence we reject the null hypothesis and confirm the discriminant validity of the scale.
@@ 5.4.3 Construct Validity
We first explore the number of factors present in the XEQ Scale using Exploratory FA. Figure 2 plots the eigenvalues for PCA coefficients derived from the scale responses. There is significant one factor that is evident throughout the scale responses which we refer to as “XAI Experience Quality”. This is evidenced by the sharp drop and plateau of eigenvalues for PCA coefficient 2 onwards.
Figure 2: Exploratory Factor Analysis
To perform the Confirmatory FA we create a factor model as suggested by exploratory FA. Confirmatory FA shows that all item factor loadings are $\geq 0.5$ baseline meaning, all items contribute to the over-arching factor measured by this scale. To validate the evaluation dimensions assigned to items by the research team, we create another factor model with 4 factors each assigned with the subset of the items indicated in Table 2 last column. Again we find that each item meets $\geq 0.5$ factor loading for their assigned factor. This confirms that while there is an over-arching factor about “XAI Experience Quality”, it is substantially underpinned by the four factors Learning, Utility, Fulfilment and Engagement. This concludes our multi-faceted refinement of the XEQ Scale based on pilot study results.
Table 2: Results
# Item I-CVI Item-Total One Factor Factor
Correlation Loading
1 The explanations received throughout the experience were consistent. 1.0000 0.6274 0.6076 Engagement
2 The experience helped me understand the reliability of the AI system. 1.0000 0.6416 0.6300 Learning
3 I am confident about using the AI system. 1.0000 0.7960 0.7790 Utility
4 The information presented during the experience was clear. 1.0000 0.7666 0.7605 Learning
5 The experience was consistent with my expectations. 0.9231 0.7959 0.7831 Fulfilment
6 The presentation of the experience was appropriate for my requirements. 0.9231 0.8192 0.8083 Fulfilment
7 The experience has improved my understanding of how the AI system works. 0.9231 0.6169 0.5859 Learning
8 The experience helped me build trust in the AI system. 0.9231 0.7160 0.7018 Learning
9 The experience helped me make more informed decisions. 0.9231 0.7460 0.7279 Utility
10 I received the explanations in a timely and efficient manner. 0.8462 0.7015 0.6841 Engagement
11 The information presented was personalised to the requirements of my role. 0.8462 0.7057 0.6801 Utility
12 The information presented was understandable within the req. of my role. 0.8462 0.7876 0.7803 Utility
13 The information presented showed me that the AI system performs well. 0.8462 0.8112 0.8016 Fulfilment
14 The experience helped to complete the intended task using the AI system. 0.8462 0.8299 0.8241 Utility
15 The experience progressed sensibly. 0.7692 0.8004 0.7912 Engagement
16 The experience was satisfying. 0.7692 0.7673 0.7529 Fulfilment
17 The information presented during the experience was sufficiently detailed. 0.7692 0.8168 0.8035 Utility
18 The experience provided answers to all of my explanation needs. 0.7692 0.8472 0.8444 Fulfilment
@@ 6.2 XEQ Scale in Practice
Table 3 presents a representative scenario of how the XEQ Scale is used in practice. It presents the XAI Experience Quality analysis of a hypothetical XAI system based on the XEQ scale administered to 5 stakeholders. The items in Table 3 are organised according to their evaluation dimensions and the stakeholder responses are randomly generated and quantified as follows: Strongly Agree-5; Somewhat Agree-4; Neutral-3; Somewhat Disagree-2; Strongly Disagree-1. Based on the responses, we calculate the following scores.
@@
quantifies the quality along each evaluation dimension and is calculated as the mean of all responses to the respective subset of items. For XAI system designers, Factor Means indicates the dimensions that need improvement. For instance, in the example XAI system, the Learning and Engagement dimensions need improvement whereas stakeholders found the Utility and Fulfilment dimensions to be satisfactory with room for improvement.
@@
quantifies the XAI experience quality of the system as a whole and is calculated as the aggregate of Factor Means. The System XEQ Score helps the XAI system designers to iteratively develop a well-rounded system grounded in user experience. The designers can also choose to assign a higher weight to one or a subset of dimensions that they find important at any iteration. System XEQ Score can also be utilised by external parties (e.g. regulators, government), either to evaluate or benchmark XAI systems.
Table 3: XEQ Scale in Practice
Factor Item# Stakeholder ID Factor
#1 #2 #3 #4 #5 Mean
Learning 2 3 2 4 3 3
4 2 1 4 3 4
7 4 3 3 4 2
8 3 3 2 1 4 2.90
Utility 3 2 4 5 3 4
9 4 2 4 4 5
11 4 2 3 5 3
12 5 2 5 2 3
14 5 3 5 3 4
17 4 2 4 4 5 3.67
Fulfilment 5 3 5 4 2 5
6 3 4 5 5 4
13 3 3 3 2 3
16 4 3 4 5 4
18 4 4 3 2 5 3.68
Engagement 1 1 2 4 2 3
10 4 3 2 4 3
15 2 1 3 3 1 2.40
Stakeholders’s XEQ 3.22 2.72 3.72 3.17 3.61
Score
System XEQ Score 3.16
@@ 6.3 XEQ Benchmark Development
The next phase for the XEQ scale entails developing a benchmark for XAI experience quality. This process includes administering the XEQ scale to over 100 real-world AI systems that provide explainability to stakeholders and establishing a classification system. We plan to follow the established benchmark maintenance policy of the User Experience Questionnaire [26] where we develop and release an XEQ Analysis tool with the benchmark updated regularly.
@@
7 Conclusion
In this paper, we presented the XEQ scale for evaluating XAI experiences. The XEQ scale provides a comprehensive evaluation for user-centred XAI experiences and fills a novel gap in the evaluation of multi-shot explanations which is currently not adequately fulfilled by any other evaluation metric(s). Throughout this paper, we have described the development and validation of the scale following psychometric theory. We make this scale available as a public resource for evaluating the quality of XAI experiences. In future work, we plan to investigate the generalisability of the XEQ scale on additional domains, AI systems and stakeholder groups. Beyond this, we propose to establish a benchmark using the XEQ scale. Our goal is to facilitate the user-centred evaluation of XAI and support the emerging development of best practices in the explainability of autonomous decision-making.
Ethical Statement
Both content validity study and pilot study protocols passed the ethics review of the leading institution (references removed for review). Informed consent was obtained from all XAI experts and pilot study participants.
{ack}
We thank all XAI experts and pilot study participants for their contributions.
This work is funded by the iSee project. iSee, an EU CHIST-ERA project, received funding for the UK from EPSRC under grant number EP/V061755/1; for Ireland from the Irish Research Council under grant number CHIST-ERA-2019-iSee and for Spain from the MCIN/AEI and European Union “NextGenerationEU/PRTR” under grant number PCI2020-120720-2.
@@ 7 Conclusion
7.1 Initial Items Bank
Table 4: Items Bank
@@ 7.1 Initial Items Bank
7.2 Content Validity Study
This study aimed to establish the content validity of the XEQ scale with XAI experts. Being XAI Experiences are a novel concept, we included three example XAI experiences that capture a variety of stakeholder types and application domains. In addition to the CourseAssist chatbot example included in the paper, they were presented with the following experiences in video format.
  • The AssistHub AI platform is a website for processing welfare applications and is used by a local council to accelerate the application process. A regulation officer is exploring the website and its XAI features to understand the fairness and bias of the AI system being used in the decision-making process. A non-interactive preview of the experience is presented in Figure 3.
@@ 7.2 Content Validity Study
Both Figures 4 and 3 are annotated with notes that describe the XAI features that were available to the stakeholder. Finally, Figure 5 presents a preview of the Study page.
@@
7.3 Pilot Study
A pilot study was conducted with 203 participants over two application domains where they evaluated either a positive or negative XAI experience. In addition to the CourseAssist chatbot examples provided in the paper, we included two XAI experiences of welfare applicants interacting with the AssistHub AI platform (see Figures 6 and 7). Notes refer to how different aspects of the explanations can lead to a positive or negative XAI experience. Similar to the previous study, all XAI experiences were available to participants in video format. Finally Figure 8 presents a preview of the Pilot study where pages 1 and 2 were customised based on the application participants were assigned to.

### v4

@@ 5.1.1 XAI experiences development and selection
Figure 2: The static previews of the relatively positive (left 2 columns) and negative (right column) XAI Experiences with the CourseAssist Chatbot designed for the Pilot Study.
5.2 Participant recruitment and selection
@@
The inclusion and exclusion criteria for participant recruitment to evaluate the CourseAssist and AssistHub experiences are summarized in Table 3. After getting through the inclusion criteria, each participant was randomly assigned to assess either the positive or negative sample experience in their respective application. After preliminary review, we excluded 5, 1 and 1 responses from groups CourseAssist-P, AssistHub-P and AssistHub-N who failed the following attention checks: 1) spend less than half of the allocated time; and/or 2) responded to the questionnaire in a pattern. This yielded 68, 70, 50, and 50 responses for the CourseAssist-P, CourseAssist-N, AssistHub-P, and AssistHub-N cohorts.
Table 3: Recruitment details for the pilot study; The only exclusion criterion was that participants of the CourseAssist study were not eligible for AssistHub, and vice versa.
Identifier Stakeholder Inclusion criteria Source Sample
CourseAssist Student Age – between 18 and 30; Current education level – Undergraduate degree; Institute 68
Degree subjects – Mathematics and statistics, Information and Communication Prolific 70
Technologies, Natural sciences; Language of instruction – English
AssistHub Welfare applicant Age – above 30; Household size – 3 or larger; Prolific 100
Property ownership – social housing or affordable-rented accommodation;
Employment status – part-time, due to start a new job within the next month,
unemployed, or not in paid work (e.g. homemaker or retired)
5.3 Data collection and analysis
The study was hosted on the Jisc Online Surveys platform. Internal consistency, discriminant validity and construct validity were evaluated with both CourseAssist and AssistHub experiences with 238 participants. Among the 68 participants from the research institute who evaluated CourseAssist experiences, 35 took part in the test-retest validity study. AssistHub experiences were excluded from the test-retest validity study due to the practical challenges of following up and re-identifying participants for retesting on the Prolific platform. The duration between the test and retest was 4 weeks.
@@ 5.3.3 Construct Validity
Construct validity evaluates the degree to which the scale assesses the characteristic of interest, i.e. factors [48]. Via factor analysis we aim to uncover underlying factors (i.e. dimensions) and validate them. Two forms of factor analysis applied are:
Exploratory Factor Analysis (EFA)
@@
Confirmatory Factor Analysis (CFA)
tests a pre-defined factor model, with factor loadings expected to be >0.5 to indicate strong support. For the XEQ scale, the evaluation dimensions assigned in Table 4 are validated as a 4-factor model.
5.3.4 Test-retest Reliability
The consistency and stability of the XEQ scale over time are measured with the test-retest reliability. We provide quantitative evidence by applying the following metrics.
Pearson Correlation Coefficient
measures the correlation between scale scores at test and retest. Let $r_{j}$ and $r_{j}^{\prime}$ represent participant $j$’s scale scores at the first instance (test) and the second instance (retest), respectively, with $\bar{r}$ and $\bar{r^{\prime}}$ being the mean scale scores for all participants at each time point. A $\rho>0.7$ indicates strong test-retest reliability, suggesting high consistency between the two measurement instances.
$\rho=\frac{\sum_{j=1}^{N}{(r_{j}-\bar{r})(r_{j}^{\prime}-\bar{r^{\prime}})}}{\sqrt{\sum_{j=1}^{N}{(r_{j}-\bar{r})^{2}}\sum_{j=1}^{N}{(r_{j}^{\prime}-\bar{r^{\prime}})^{2}}}}$
Intra-class Correlation Coefficient (ICC)
assesses the consistency of scale scores from the same group of participants between test and retest, using a two-way mixed-effects model (i.e. ICC3). Given that each participant $j$ has a mean score of $\bar{r_{j}}$ between test and retest, and the overall mean across all participants is $\overline{\overline{r}}$ the ICC is computed as:
$ICC=\frac{ms-ms_{e}}{ms+(k-1)\times ms_{e}}$
, where $ms$ is the mean square between participants, $ms=\sum_{j=1}^{N}{(\bar{r_{j}}-\overline{\overline{r}})^{2}}/(N-1)$ and $ms_{e}$ is the mean square error $ms_{e}=\sum_{j=1}^{N}{[(r_{j}-\bar{r_{j}})^{2}+(r_{j}^{\prime}-\bar{r_{j}})^{2}]}/[N\times(k-1)]$. An ICC value above 0.75 indicates good reliability, and above 0.90 reflects excellent reliability, indicating strong consistency in scale scores between test and retest.
Table 4: Results: 18 items were selected from the item bank based on the Item-Level Content Validity Index(I-CVI> 0.75) and were assigned an evaluation dimension. From the pilot studies, we find Item-Total Correlation (iT)>0.5 for all 18 items. One-factor loadings all over 0.5 confirm the overarching single factor within the XEQ Scale. Confirmatory (C.) Factor Analysis shows significant loading (>0.5) from each item for their respective evaluation dimension.
# Item I-CVI $iT$ One-Factor Evaluation C. Factor
Loading Dimension Loading
1 The explanations received throughout the experience were consistent. 1.0000 0.6274 0.6076 Engagement 0.7988
2 The experience helped me understand the reliability of the AI system. 1.0000 0.6416 0.6300 Learning 0.9409
3 I am confident about using the AI system. 1.0000 0.7960 0.7790 Utility 1.0096
4 The information presented during the experience was clear. 1.0000 0.7666 0.7605 Learning 1.1634
5 The experience was consistent with my expectations. 0.9231 0.7959 0.7831 Fulfilment 1.0822
6 The presentation of the experience was appropriate for my requirements. 0.9231 0.8192 0.8083 Fulfilment 0.9300
7 The experience has improved my understanding of how the AI system works. 0.9231 0.6169 0.5859 Learning 0.8280
8 The experience helped me build trust in the AI system. 0.9231 0.7160 0.7018 Learning 0.9787
9 The experience helped me make more informed decisions. 0.9231 0.7460 0.7279 Utility 0.7909
10 I received the explanations in a timely and efficient manner. 0.8462 0.7015 0.6841 Engagement 0.8592
11 The information presented was personalised to the requirements of my role. 0.8462 0.7057 0.6801 Utility 0.7938
12 The information presented was understandable within the requirements of my role. 0.8462 0.7876 0.7803 Utility 1.1452
13 The information presented showed me that the AI system performs well. 0.8462 0.8112 0.8016 Fulfilment 0.9071
14 The experience helped to complete the intended task using the AI system. 0.8462 0.8299 0.8241 Utility 1.0787
15 The experience progressed sensibly. 0.7692 0.8004 0.7912 Engagement 1.1588
16 The experience was satisfying. 0.7692 0.7673 0.7529 Fulfilment 1.0309
17 The information presented during the experience was sufficiently detailed. 0.7692 0.8168 0.8035 Utility 1.0299
18 The experience provided answers to all of my explanation needs. 0.7692 0.8472 0.8444 Fulfilment 0.9041
@@ 5.4.1 Internal Consistency
Table 4 column $iT$ reports the Item-Total Correlation. All items met the baseline criteria of $iT\geq 0.5$ and baseline criteria for Inter-Item correlation. Cronbach’s alpha is $0.9562$ which also indicates strong internal consistency.
@@ 5.4.2 Discriminant Validity
We performed discriminant analysis over 100 trials where at each trial a different train-test split of the responses was used. Each trial used a stratified split, with 70% of the responses for training and 30% for testing. Over the 100 trials, we observed accuracy of $0.63\pm 0.05$ and a macro F1-score of $0.63\pm 0.05$ which is significantly over the baseline accuracy of $0.50$ for a binary classification task. Mixed-model ANOVA test showed a statistically significant difference between groups P and N with a p-value of $121.63e-$ 12 where the mean participant total for groups P and N were $70.96\pm 0.47$ and $57.97\pm 1.84$. Also, it revealed a substantial variability within groups indicated by the group variance of 104.86, which we account for including responses from two application domains. Furthermore, Cohen’s d was 1.7639 which indicates a large effect size confirming a significant difference between groups A and B. A standard t-test also obtained a p-value of $091.13e-$ 09 further verifying the statistically significant difference. Based on this evidence we reject the null hypothesis and confirm the discriminant validity of the scale.
@@ 5.4.3 Construct Validity
We first explore the number underlying factors in the XEQ Scale using EFA. Figure 3 presents the eigenvalues for PCA coefficients derived from scale responses which shows a sharp drop and plateau of eigenvalues for coefficient 2 onwards. This signifies a single overarching factor evident throughout the scale which we refer to as “XAI Experience Quality”. We further confirm this single factor via one-factor loadings (i.e. CFA with a single factor model) given in Table 4 where all item supports are > 0.5.
Figure 3: Scree plot of eigenvalues derived from the Principal Component Analysis (PCA) performed for the exploratory factor analysis (EFA); a sharp decline in eigenvalues indicates the presence of a single overarching factor.
An additional CFA was conducted to confirm that evaluation dimensions had substantial factor loadings from their respective items. The rightmost column in Table 4 presents the CFA factor loadings. Each item meets the required factor loading of $\geq 0.5$ for their evaluation dimension. These observations reinforce that while there is an over-arching factor on “XAI Experience Quality”, it is substantially underpinned by the four evaluation dimensions Learning, Utility, Fulfilment and Engagement.
5.4.4 Test-retest Reliability
The test-retest reliability of the scale was assessed with a 4-week interval using the two CourseAssist experiences. Pearson correlation coefficient $\rho=0.7561$ with a p-value of $071.5e-$ 07, indicated a strong linear relationship between the test and retest scores. Additionally, $ICC=0.8609$ with a p-value of $083.17e-$ 08, falls within the range that indicates strong consistency. Both measures establish the test-retest reliability of the XEQ scale.
This concludes our multi-faceted evaluation and refinement of the XEQ Scale based on pilot study results.
@@ 6.2 XEQ scale in practice
We propose the XEQ Scale as a tool to support the development and evaluation of interactive user-centric XAI systems. We anticipate calculating 3 aggregate metrics based upon responses to the 18 items in the scale:
@@
quantifies the quality along each evaluation dimension and is calculated as the mean of all responses to the respective subset of items. For XAI system designers, a lower factor score indicates the dimensions that need improvement.
@@
quantifies the XAI experience quality of the system as a whole and is calculated as the aggregate of Factor scores. The System XEQ Score helps the XAI system designers to iteratively develop a well-rounded system grounded in user experience. The designers can also choose to assign a higher weight to one or a subset of dimensions that they find important at any iteration. System XEQ Score can also be utilised by external parties (e.g. regulators, government), either to evaluate or benchmark XAI systems.
We have created the following user story to demonstrate the usefulness of the XEQ Scale for the iterative development and evaluation of an applied XAI system.
Jane is the systems manager for a motor insurance company. The company have recently introduced an AI system for calculating an insurance quote using applicants’ immediate driving history and previous claims. As part of the company’s regulatory obligations, the AI system is supported by several explanation algorithms. However, the systems management team are having complaints from sales staff that they struggle to justify system outcomes using these explanations. Jane therefore decides to use the XEQ Scale to measure the XE for 2 stakeholder groups: customers and sales staff. They start by recruiting a representative user group for each role. Subsequently, each participant used the XAI system and scored their experience with the XEQ Scale.
Jane collates the responses and calculates the stakeholder scores, factor scores and system scores for each role. When comparing system scores, Jane finds that sales staff consistently score the XAI system lower compared to customers. Looking at granular factor scores, they identify that the Utility factor score is particularly low for sales staff, specifically, item 14 (’The experience helped me to complete the intended task using the AI system’) received consistently negative feedback. From these insights, Jane recognises that an additional explanation algorithm should be added to support sales staff in task completion. Jane updated the explanation algorithms and trials once again with a sample of sales staff, finding that the new system increased the Utility factor score and overall, led to an improved system XEQ score.
From the above user story, it can be observed that the XEQ Scale can support the iterative development of XAI systems by facilitating the targeted update via user-centred evaluation.
@@ 6.3 XEQ Benchmark Development
The next phase for the XEQ scale entails developing a benchmark for XAI experience quality. This process includes administering the XEQ scale to over 100 real-world AI systems that provide explainability to stakeholders and establishing a classification system. We are currently following the established benchmark maintenance policy of the User Experience Questionnaire [49] where we develop and release an XEQ Analysis tool with the benchmark updated regularly.
@@
7 Conclusions
In this paper, we presented the XEQ scale. The XEQ scale provides a framework for the comprehensive evaluation of user-centred XAI experiences. It fills a novel gap in the evaluation of multi-shot explanations which is currently not adequately fulfilled by any other evaluation metric(s). Throughout this paper, we have described the development and validation of the scale following psychometric theory. We make this scale available as a public resource for evaluating the quality of XAI experiences. In future work, we plan to investigate the generalisability of the XEQ scale on additional domains, AI systems and stakeholder groups. Beyond this, we propose to establish a benchmark using the XEQ scale. Our goal is to facilitate the user-centred evaluation of XAI and support the emerging development of best practices in the explainability of autonomous decision-making.
Acknowledgements.
The authors thank all participants in the CVR and pilot studies. iSee is an EU CHIST-ERA project that received funding for the UK from EPSRC under grant number EP/V061755/1, for Ireland from the Irish Research Council under grant number CHIST-ERA-2019-iSee, for France from the French National Research Agency under grant number ANR-21-CHR4-0004, and for Spain from the MCIN/AEI and European Union “NextGenerationEU/PRTR” under grant number PCI2020-120720-2.
@@ Acknowledgements.
A: Initial Items Bank
Table 5 presents the initial item bank compiled from literature and reviewed by the research team.
Table 5: Initial Items Bank
@@
B: Content Validity Study
This study aimed to establish the content validity of the XEQ scale with XAI experts. Since XAI Experiences are a novel concept, we included three example XAI experiences that capture a variety of stakeholder types and application domains. In addition to the CourseAssist chatbot example included in the paper, they were presented with the following experiences in video format.
  • The AssistHub AI platform is a website for processing welfare applications and is used by a local council to accelerate the application process. An auditor is exploring the website and its XAI features to understand the fairness and bias of the AI system being used in the decision-making process. A non-interactive preview of the experience is presented in Figure 4.
@@
Both Figures 5 and 4 are annotated with notes that describe the XAI features that were available to the stakeholder. Finally, Figure 6 presents a preview of the Study page.
@@
C: Pilot Study
A pilot study was conducted with 238 participants over two application domains where they evaluated either a positive or negative XAI experience. In addition to the CourseAssist chatbot examples provided in the paper, we included two XAI experiences of welfare applicants interacting with the AssistHub AI platform (see Figures 7 and 8). Notes refer to how different aspects of the explanations can lead to a positive or negative XAI experience. Similar to the previous study, all XAI experiences were available to participants in video format. Finally Figure 9 presents a preview of the Pilot study where pages 1 and 2 were customised based on the application participants were assigned to.